OpenStudy (anonymous):

FAN AND MEDAL!!!!! Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). A) y = -1/3x - 6 B) y = -1/3x - 14 C) y = 3x - 6 D) y = 3x - 14

OpenStudy (anonymous):

I think its C but I'm not sure. can anyone check my answer?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

the y = 3x is correct, lets check the constant

OpenStudy (anonymous):

@kai_is_my_name You can use this to graph: https://www.desmos.com/calculator

OpenStudy (amistre64):

-2 = 3(4) + c -2 = 12 + c -14 = c

OpenStudy (amistre64):

looks like -6 aint gonna cut it

OpenStudy (anonymous):

Just plug in 3x - y = -3, (4, -2), y = -1/3x - 6, y = -1/3x - 14, y = 3x - 6, and y = 3x - 14 into separate boxes on the left.

OpenStudy (anonymous):

Then find out which one is parallel and crosses (4, -2)

OpenStudy (anonymous):

ok, thank you. @amistre64 so what your saying is that the correct y- intercept is -14 so its D?

OpenStudy (amistre64):

correct :)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

@kai_is_my_name That is correct.

OpenStudy (amistre64):

youre welcome, and good luck :)

OpenStudy (anonymous):

Parallel lines = same slope.